# Identifying the Desired Word Suggestion in Simultaneous Audio

### v1

@@ Abstract.
ccs: Human-centered computing Accessibility technologiesccs: Human-centered computing Sound-based input / outputccs: Human-centered computing Text input
@@ 1. Introduction
Entering text on a device such as a smartphone is a common daily task. Commercial touchscreen keyboards often place a number of word suggestions, commonly three or four, above the keyboard. This allows a user to quickly glance and check the suggestions without distracting too much from the task of typing. If a user spots their desired word, they can tap it to write it and avoid typing the entire word letter-by-letter. However, for users with visual impairments, it can be a challenge to utilize word suggestions that are presented in such a visual manner. Screen reading software such as Apple’s VoiceOver or Google’s TalkBalk requires a user touch these word suggestions or perform a special gesture to hear them read, which can interrupt the flow of typing.
In text entry interfaces, suggested words and word completions could be offered if certain conditions are met, such as likelihood of the suggestion surpassing a threshold (Montague et al., 2016; Nicolau et al., 2019) or the user typing a certain number of characters (Kristensson and Müllners, 2021). For non-visual interfaces, audio suggestions could be provided automatically when these conditions occur instead of requiring the user to take an action to trigger them. If the user hears the word they are typing they could select it immediately, without needing to finish the remainder of the letters. This would operate similar to many commercially available standard touchscreen keyboards, but using audio instead of visual feedback.
Reading three or four word suggestions sequentially using text-to-speech (TTS) can take quite some time even with accelerated TTS. The user would likely be able to type the remainder of the word by the time the list was read. To present suggested words more efficiently, we investigate providing the user with audio-only word suggestions using simultaneous voices as suggested by Montague et al. (Montague et al., 2016). This would enable the interface to provide word suggestions much quicker and potentially increase the utility of word predictions for eyes-free text input. In this work, we conduct two perceptual studies to measure user performance at detecting a target word among several potentially similar sounding words (as is often the case for a set of word suggestions). We aim to determine not only if users can detect the presence of their target word, but also if they can identify which voice said it as might be required to select their target word in a text entry interface.
@@
  • We conduct the first quantitative study into if users can find a target word within a set of similar words played via simultaneous TTS. The word sets we tested were based on realistic word suggestions made in the context of typing real-world sentences in conjunction with a word prediction algorithm.
@@ 2. Related Work
There have been several studies on the human ability to perceive speech amid other speech, the first by Cherry in 1953 (Cherry, 1953). Cherry conducted a series of experiments on the Cocktail Party Effect. Cherry found participants could clearly listen to either of two spoken messages recorded by the same person played into opposite ears via headphones. However, participants were not able to notice when the language of the message in the ear that they were not listening to (the “rejected” ear) changed from English to German. This prompted further studies, in which the Cherry found that participants were able to correctly identify if the rejected message was normal human speech (as opposed to reversed speech or an oscillating tone), as well as whether it was a male or female voice. They were not, however, able to identify any words in the rejected message or what language it was in.
This line of work was continued by Egan et al. (Egan et al., 1954), who played sentences simultaneously for participants. Each of these sentences contained one of two “call signs”. Participants were instructed to write down the words following a particular call sign, ignoring the sentence that contained the other call sign. In one experiment, they found that high-pass filtering either message allowed participants to better understand the target message. This suggests that users will be able to understand simultaneous speech better if there are differences in the sound frequencies. In another experiment, Egan et al. confirmed the results found by Cherry (Cherry, 1953) that participants were better able to understand the target message when the two messages were played in opposite ears instead of the same ear.
@@
Brungart and Simpson (Brungart and Simpson, 2007) experimented with two, three, and four speakers in different spatial arrangements around the listener. They used digital signal processing and head-related transfer functions to “move” the stimuli in space around a user’s head. They used the same phrase structure (call sign, color, number) as Brungart (Brungart, 2001). Their results showed that participants were 92% accurate with two speakers, 72% with three speakers, and 62% with four. Additionally, they found that participants scored higher in spatial configurations where the speakers were more spread out, and specifically when the target phrase came from either the left or right side as opposed to the middle.
@@ 3. Study 1
Prior to the study, we created the trials based on phrases taken from the Enron mobile data set (Vertanen and Kristensson, 2011). To create the words spoken for each trial, we chose a random Enron phrase and a random character position in that phrase. We fed the portion of the phrase up to and including that position (including any partial word) into the VelociTap decoder (Vertanen et al., 2015). The n-best word suggestions returned by the decoder (the $n$ words with the highest probability given the context and any prefix) became the words spoken by each voice, where $n$ was the number of voices in that particular trial. We did not allow the random position to be the beginning of the sentence since the decoder would just return the most likely words to start a sentence. If there were not enough words returned by the decoder to fill every voice in a trial, a new phrase and position were chosen. This procedure produced word suggestions sets that were representative of what a user might receive during real-world text entry. The partially completed word or next word (if the position fell between words) became the target word. Once the words were selected, which voices spoke them were randomized so that Voice 1 was not always the most likely word returned by the decoder. It is important to note that the target word was not always present in the words that were spoken, as would be the case in real-world text entry (this was the case in 22% of trials).
@@ 3.1. Procedure
We developed a web application to conduct this study remotely. This allowed us to ensure participant safety during the COVID-19 pandemic. The application first walked participants through informed consent and a demographic questionnaire. It then provided instructions for the first of the three counterbalanced conditions; the ordering of conditions was predetermined based on participant number. To familiarize participants with the voices and their locations for each condition, we included a button for each voice on the instruction page, as shown in Figure 1. When clicked, these buttons played audio speaking their voice number (e.g. Voice 1 played the text-to-speech for “Voice 1”). We synthesized the voice number using the corresponding voice, and spatially located each voice in its position for the current condition. We required participants click all four buttons before they could proceed. They were allowed to listen to each voice as many times as they wanted. The instructions also requested that participants wear headphones for the study. Participants completed this instruction process at the start of each condition.
Figure 1. The web interface at the beginning of the Sequential condition. The text below the Voice buttons as well as the Continue button do not appear until all four Voice buttons have been clicked. This screen is meant to familiarize users with which voice corresponds to which number.
@@
The instructions read, “In the following study you will be asked to identify which voice spoke a specified word. There will be up to four voices in various spatial arrangements. There are a total of 3 conditions. If you have headphones available, please use them for this study. In this condition, the voices will be played sequentially. Voice 1 (male) will be played first, followed by Voice 2 (female), Voice 3 (male), and then Voice 4 (female). Select each button below to hear each voice separately.” Below the button for each voice, the instructions continue, “On the next page, the target word will be displayed. When you are ready, please select the Ready button and the audio will play. You will then be able to select which voice you heard speak the target word. The first 3 trials will be practice to get you aquainted with the interface. Please select continue.”
Practice trials were not included in any of the data analysis. Once the practice trials were complete, participants then completed 24 evaluation trials. The interface was identical to that used in the practice trials, but participants were not able to retry any incorrect trials. At the end of each condition, participants completed a brief questionnaire. They then repeated this process for the next condition, beginning with the instructions page.
Figure 2. The web interface once the participant has clicked ‘Ready’ and the audio has played. The participant now selects which voice they heard say the target word.
@@
3.2. Results
@@
There were two independent variables in this study: spatial arrangement and number of voices. The dependent variable was response accuracy, which we computed as the number of correct responses divided by the number of total responses. If a participant selected “I don’t know”, this was considered an incorrect response. The mean response accuracy for each combination of spatial arrangement and number of voices can be seen in Figure 3. For the 180Degree condition, participants had an average accuracy of 85%, 65%, 41% with two, three, and four voices, respectively. The results in the 60Degree condition were similar, with an accuracy of 84% for two voices, 61% for three voices, and 43% for four voices. As we expected, the Sequential condition did not have the same decrease in accuracy when more voices were added. Participants had accuracies of 96% (two voices), 92% (three voices), and 92% (four voices).
[…]
We ran pairwise post hoc t-tests with Bonferroni correction to determine where exactly the significant differences were found. The results of these tests are shown in Table 1. Notably, within both 180Degree and 60Degree the difference between each number of voices was significant. Trials that had three voices had significantly lower accuracy than trials with two voices, and trials with four voices were significantly lower still. However, within Sequential, the number of voices did not significantly impact accuracy. Between the two simultaneous conditions, 180Degree and 60Degree, differences were significant only where the number of voices was not equal. This suggests that there was not a large difference between these two spatial arrangements. Finally, the difference between each simultaneous condition with two voices and Sequential with any number of voices was not significant, suggesting that simultaneous audio could be used if there were only two word suggestions. However, we do not recommend using simultaneous audio for more than two words since three and four voices within the simultaneous conditions were significantly worse than Sequential with any number of voices.
@@ 3.3. Discussion
This first study showed that performance did not significantly change between the conditions when there were only two voices. However, participant accuracy was significantly lower in both simultaneous conditions when there were more than two voices, compared to both when there were only two voices and to the Sequential control condition. We were hoping to have success similar to past work by creating both spatial and pitch differences in our speakers. However, we conjecture that the difficulty for participants arose from the short length and similarity of the stimuli. All of the previous studies we discussed earlier in this work used sentence or phrase stimuli that may have allowed participants more time to adjust to hearing multiple speakers. Additionally, many previous studies, like those by Brungart (Brungart, 2001) and Brungart and Simpson (Brungart and Simpson, 2007), used a limited set of words in their phrases that were not easily confusable. This was not the case in our study, and that may have also contributed to the poorer performance observed in the simultaneous conditions. We also reduced the likelihood of participants guessing the correct voice by instructing them to select “I don’t know” if they were unsure, since it is likely that a user entering text would not elect to select a word suggestion if they were not sure which to select.
@@ 4.1. Procedure
This study used the same web interface as Study 1. Participants completed informed consent, a demographic questionnaire and received instructions before completing six practice trials and 24 evaluation trials in each condition. There was a short questionnaire after each condition. To recruit a larger and more diverse set of participants in this study, we recruited 96 participants by posting Human Intelligence Tasks (HITs) on Amazon Mechanical Turk, a crowdsourcing platform. We paid participants \$3.50 for their time. Similar to the crowdsourcing procedure used by Vertanen et al. (Vertanen et al., 2018), we eliminated workers that failed to meet a control standard. The standard we chose was 50% accuracy in the Sequential (control) condition. Since there was no overlap between the voices in the Sequential condition, this should have been an easy accuracy to achieve. We returned rejected HITs to the pool until we had 96 participants that met the standard. No participant was able to complete more than one HIT in this study, and, to our knowledge, no participants had previously completed Study 1 since we recruited from different pools for each study.
@@ 4.2. Results
In this study, our independent variables were the length of the delay and the number of voices. As with the previous study, our dependent variable was accuracy. The mean accuracy results broken down by each combination of delay and number of voices can be seen in Figure 4. In the DelayShort condition, participants had average response accuracies of 84% with two voices, 69% with three voices, and 60% with four voices. Increasing the delay to 0.25 s resulted in accuracies of 88%, 79%, and 76% with two, three, and four voices, respectively. The baseline Sequential condition yielded accuracies of 86% (two voices), 89% (three voices), and 90% (four voices). It was interesting to note that accuracy increased with the number of voices in the Sequential condition. One possible explanation for this is that participants may have selected Voice 2, for example, if the second voice spoke their word, even if only Voices 1 and 4 were speaking, as was the case when only two voices were present. Participants selected a voice that was not a valid option infrequently, with this occurring in only 6.4% of trials with less than four voices across all conditions. However, we did notice this was the highest in Sequential with two voices at 10.5% of trials. Further clarification in the instructions or disabling the buttons for voices that were not a part of a particular trial may reduce this occurrence in future work.
[…]
We ran the pairwise post hoc t-tests with Bonferroni correction to locate the significant differences. These results are presented in Table 2. While, as we noted, mean accuracy increased slightly with number of voices in the Sequential condition, none of these differences were significant. In the DelayLong condition, participants’ accuracy with two voices was significantly higher than with three or four, but there was no significant difference between three and four voices. In the DelayShort condition, accuracy significantly decreased with each voice added.
@@
Table 3. The average playing time in seconds for the trials in each condition and for different number of simultaneous voices.
4.3. Discussion and Limitations
Adding the slight delay before starting to play each subsequent voice produced much more favorable results than the spatial and pitch distance alone, especially for greater numbers of voices. While it is difficult to directly compare the DelayShort and DelayLong conditions to the 60Degree condition from Study 1 given the different participant pools, the mean accuracies were much higher: 60.2% and 76.0% with four voices in DelayShort and DelayLong, respectively, compared to 42.7% in 60Degree from Study 1.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: The PErvasive Technologies Related to Assistive Environments; June 25–27, 2025; Corfu Island, Greecebooktitle: The PErvasive Technologies Related to Assistive Environments (PETRA ’25), June 25–27, 2025, Corfu Island, Greecedoi: 10.1145/3733155.3733212isbn: 979-8-4007-1402-3/25/06ccs: Human-centered computing Accessibility technologiesccs: Human-centered computing Sound-based input / outputccs: Human-centered computing Text input
@@ 1. Introduction
Entering text on a device such as a smartphone is a common daily task. Commercial touchscreen keyboards often place a number of word suggestions, commonly three or four, above the keyboard. This allows a user to quickly glance above the keyboard to check the suggestions without unduly distracting from the task of typing. If a user spots their desired word, they can tap it to write it and avoid typing the entire word letter-by-letter. However, for users with visual impairments, it can be a challenge to utilize word suggestions that are presented in such a visual manner. Screen reading software such as Apple’s VoiceOver or Google’s TalkBack requires a user touch these word suggestions or perform a special gesture to hear them read. This makes the suggestions difficult to discover and can interrupt the natural flow of typing.
A text entry interface can offer word suggestions when certain conditions are met, such as when the likelihood of a suggestion surpasses a threshold (Montague et al., 2016; Nicolau et al., 2019) or after the user types a certain number of characters (Kristensson and Müllners, 2021). For non-visual interfaces, audio suggestions could be provided automatically when these conditions occur instead of requiring the user to take an action to trigger them. If the user hears the word they are typing they could select it immediately, without needing to finish the remainder of the letters. This would operate similar to many commercially available standard touchscreen keyboards, but using audio instead of visual feedback.
Reading three or four word suggestions sequentially using text-to-speech (TTS) can take quite some time even with accelerated TTS. The user would likely be able to type the remainder of the word by the time the list was read. To present suggested words more efficiently, we investigate providing the user with audio-only word suggestions using simultaneous voices as suggested by Montague et al. (Montague et al., 2016). This would enable the interface to provide word suggestions much quicker and potentially increase the utility of word predictions for eyes-free text input.
In this work, we conduct two perceptual studies to measure user performance at detecting a target word among several potentially similar sounding words (as is often the case for a set of word suggestions). We aim to determine not only if users can detect the presence of their target word, but also if they can identify which voice said it as might be required to select their target word in a text entry interface.
@@
  • We conduct the first quantitative study into whether users can find a target word within a set of similar words played via simultaneous TTS. Our word sets were based on realistic word suggestions made in the context of typing real-world sentences in conjunction with a word prediction algorithm.
@@ 2. Related Work
There have been several studies on the human ability to perceive speech amid other speech, the first by Cherry in 1953 (Cherry, 1953). Cherry conducted a series of experiments on the Cocktail Party Effect. Cherry found participants could clearly listen to either of two spoken messages recorded by the same person played into opposite ears via headphones. However, participants were not able to notice when the language of the message in the ear that they were not listening to (the “rejected” ear) changed from English to German. This prompted further studies, in which the Cherry found that participants were able to correctly identify if the rejected message was normal human speech (as opposed to reversed speech or an oscillating tone), as well as whether it was a male or female voice. However, they were not able to identify any words in the rejected message or what language it was in.
This line of work was continued by Egan et al. (Egan et al., 1954) who had participants listen to sentences played simultaneously. Each sentence contained one of two “call signs”. Participants were instructed to write down the words following a particular call sign, ignoring the sentence that contained the other call sign. In one experiment, they found that high-pass filtering either message allowed participants to better understand the target message. This suggests that users will be able to understand simultaneous speech better if the voices have different sound frequencies. In another experiment, Egan et al. confirmed the results found by Cherry (Cherry, 1953) that participants could understand the target message better if the two messages were played in opposite ears instead of the same ear.
@@
Brungart and Simpson (Brungart and Simpson, 2007) experimented with two, three, and four speakers in different spatial arrangements around the listener. They used digital signal processing and head-related transfer functions to “move” the stimuli in space around a user’s head. They used the same phrase structure (call sign, color, number) as Brungart (Brungart, 2001). Their results showed that participants were 92% accurate with two speakers, 72% with three speakers, and 62% with four speakers. Additionally, they found that participants scored higher in spatial configurations where the speakers were more spread out, and specifically when the target phrase came from either the left or right side as opposed to the middle.
@@ 3. Study 1
Prior to the study, we created the trials based on phrases taken from the Enron mobile data set (Vertanen and Kristensson, 2011). To create the words spoken for each trial, we chose a random Enron phrase and a random character position in that phrase. We fed the portion of the phrase up to and including that position (including any partial word) into the VelociTap decoder (Vertanen et al., 2015). The n-best word suggestions returned by the decoder (the $n$ words with the highest probability given the context and any prefix) became the words spoken by each voice, where $n$ was the number of voices in that particular trial.
We did not allow the random position to be the beginning of the sentence since the decoder would just return the most likely words to start a sentence. If there were not enough words returned by the decoder to fill every voice in a trial, a new phrase and position were chosen. This procedure produced word suggestions sets that were representative of what a user might receive during real-world text entry. The partially completed word or next word (if the position fell between words) became the target word. Once the words were selected, which voices spoke them were randomized so that Voice 1 was not always the most likely word returned by the decoder. It is important to note that the target word was not always present in the words that were spoken, as would be the case in real-world text entry (this was the case in 22% of trials).
@@ 3.1. Procedure
We conducted this study remotely via a web application. This allowed us to ensure participant safety during the COVID-19 pandemic. The application first walked participants through informed consent and a demographic questionnaire. It then provided instructions for the first of the three counterbalanced conditions; the ordering of conditions was determined based on participant number.
To familiarize participants with the voices and their locations for each condition, we included a button for each voice on the instruction page (Figure 1). When clicked, these buttons played audio speaking their voice number (e.g. Voice 1 played the text-to-speech for “Voice 1”). We synthesized the voice number using the corresponding voice and spatially located each voice in its position for the current condition. We required participants click all four buttons before they could proceed. They were allowed to listen to each voice as many times as they wanted. The instructions also asked participants to wear headphones for the study. Participants completed this instruction process at the start of each condition.
Figure 1. The web interface at the beginning of the Sequential condition. The text below the Voice buttons as well as the Continue button did not appear until all four Voice buttons have been clicked. This screen is meant to familiarize users with which voice corresponds to which number.
@@
The instructions read, “In the following study you will be asked to identify which voice spoke a specified word. There will be up to four voices in various spatial arrangements. There are a total of 3 conditions. If you have headphones available, please use them for this study. In this condition, the voices will be played sequentially. Voice 1 (male) will be played first, followed by Voice 2 (female), Voice 3 (male), and then Voice 4 (female). Select each button below to hear each voice separately.” Below the button for each voice, the instructions continue, “On the next page, the target word will be displayed. When you are ready, please select the Ready button and the audio will play. You will then be able to select which voice you heard speak the target word. The first 3 trials will be practice to get you acquainted with the interface. Please select continue.”
Figure 2. The web interface once the participant had clicked ‘Ready’ and the audio had played. The participant now selects which voice they heard say the target word.
@@
Practice trials were excluded from our analysis. After the practice trials, participants completed 24 evaluation trials. The interface was identical except participants were unable to retry incorrect trials. At the end of each condition, participants completed a brief questionnaire. They then repeated this process for the next condition, beginning with the instructions page.
@@
3.2. Results
@@
Our study had two independent variables: spatial arrangement and number of voices. The dependent variable was response accuracy, which we computed as the number of correct responses divided by the number of total responses. If a participant selected “I don’t know”, this was considered an incorrect response.
The mean response accuracy for each combination of spatial arrangement and number of voices can be seen in Figure 3. For the 180Degree condition, participants had an average accuracy of 85%, 65%, 41% with two, three, and four voices, respectively. The results in the 60Degree condition were similar, with an accuracy of 84% for two voices, 61% for three voices, and 43% for four voices. As we expected, the Sequential condition did not have the same decrease in accuracy as voices were added. Participants had accuracies of 96% (two voices), 92% (three voices), and 92% (four voices).
[…]
We ran pairwise post hoc t-tests with Bonferroni correction to locate any significant differences. The results of these tests are shown in Table 1. Notably, within both 180Degree and 60Degree, the difference between each number of voices was significant. Trials that had three voices had significantly lower accuracy than trials with two voices, and trials with four voices were significantly lower still. However, within Sequential, the number of voices did not significantly impact accuracy.
Between the two simultaneous conditions, 180Degree and 60Degree, differences were significant only where the number of voices was not equal. This suggests that there was not a large difference between these two spatial arrangements.
Finally, the difference between each simultaneous condition with two voices and Sequential with any number of voices was not significant, suggesting that simultaneous audio could be used if there were only two word suggestions. However, we do not recommend using simultaneous audio for more than two words since three and four voices within the simultaneous conditions were significantly worse than Sequential with any number of voices.
@@ 3.3. Discussion
Our first study showed that performance did not significantly change when there were only two voices. However, participant accuracy was significantly lower in both simultaneous conditions when there were more than two voices. This was true both when there were two voices and in the Sequential control condition.
We were hoping to have success similar to past work by creating both spatial and pitch differences in our speakers. However, we conjecture that the difficulty for participants arose from the short length and similarity of the stimuli. All of the previous studies we discussed earlier in this work used sentence or phrase stimuli that may have allowed participants more time to adjust to hearing multiple speakers.
Additionally, many previous studies, like those by Brungart (Brungart, 2001) and Brungart and Simpson (Brungart and Simpson, 2007), used a limited set of words in their phrases that were not easily confusable. This was not the case in our study, and that may have also contributed to the poorer performance observed in the simultaneous conditions. We also reduced the likelihood of participants guessing the correct voice by instructing them to select “I don’t know” if they were unsure, since it is likely that a user entering text would not elect to select a word suggestion if they were not sure which to select. Previous studies (e.g. (Brungart and Simpson, 2007; Brungart, 2001) did not report having an “I don’t know” option available.
@@ 4.1. Procedure
This study used the same web interface as Study 1. Participants completed informed consent, a demographic questionnaire and received instructions before completing six practice trials and 24 evaluation trials in each condition. There was a short questionnaire after each condition. To recruit a larger and more diverse set of participants in this study, we recruited 96 participants by posting Human Intelligence Tasks (HITs) on Amazon Mechanical Turk, a crowdsourcing platform. We paid participants \$3.50 for their time.
Similar to the crowdsourcing procedure used by Vertanen et al. (Vertanen et al., 2018), we eliminated workers that failed to meet a control standard. The standard we chose was 50% accuracy in the Sequential (control) condition. Since there was no overlap between the voices in the Sequential condition, this should have been an easy accuracy to achieve. We returned rejected HITs to the pool until we had 96 participants that met the standard. No participant was able to complete more than one HIT in this study, and, to our knowledge, no participants had previously completed Study 1 since we recruited via word-of-mouth in Study 1 and not via Amazon Mechanical Turk.
@@ 4.2. Results
In this study, our independent variables were the length of the delay and the number of voices. As with Study 1, our dependent variable was accuracy. The mean accuracy results broken down by each combination of delay and number of voices can be seen in Figure 4. In the DelayShort condition, participants had average response accuracies of 84% with two voices, 69% with three voices, and 60% with four voices. Increasing the delay to 0.25 s resulted in accuracies of 88%, 79%, and 76% with two, three, and four voices, respectively. The baseline Sequential condition yielded accuracies of 86% (two voices), 89% (three voices), and 90% (four voices).
It was interesting to note that accuracy increased with the number of voices in the Sequential condition. One possible explanation for this is that participants may have selected Voice 2, for example, if the second voice spoke their word, even if only Voices 1 and 4 were speaking, as was the case when only two voices were present. Participants selected a voice that was not a valid option infrequently, with this occurring in only 6.4% of trials with less than four voices across all conditions. However, we did notice this was the highest in Sequential with two voices at 10.5% of trials. Further clarification in the instructions or disabling the buttons for voices that were not a part of a particular trial may reduce this occurrence in future work.
[…]
We ran pairwise post hoc t-tests with Bonferroni correction to locate any significant differences. These results are presented in Table 2. While, as we noted, mean accuracy increased slightly with number of voices in the Sequential condition, none of these differences were significant. In the DelayLong condition, participants’ accuracy with two voices was significantly higher than with three or four, but there was no significant difference between three and four voices. In the DelayShort condition, accuracy significantly decreased with each voice added.
@@
Table 3. The average playing time in seconds for each condition and for different number of simultaneous voices.
4.3. Discussion
Adding a slight delay before starting to play each subsequent voice produced much more favorable results than the spatial and pitch distance alone, especially for a greater number of voices. While it is difficult to directly compare the DelayShort and DelayLong conditions to the 60Degree condition from Study 1 given the different participant pools, the mean accuracies were much higher: 60.2% and 76.0% with four voices in DelayShort and DelayLong, respectively, compared to 42.7% in 60Degree from Study 1.